OpenStudy (anonymous):

I need help!!!!!!!! An airplane flew 3 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air. -180 miles? -200 miles? -220 miles? -240 miles?

OpenStudy (anonymous):

Let v be the speed of the plane. t1=3 hrs t2=2 hrs

OpenStudy (anonymous):

(v-40) t1= s miles (v+40) t2=s miles Now equate the two equations, substitute for t1 & t2 & find v.

OpenStudy (anonymous):

(v-40) 3 = (v+40) 2 3v-120=2v+80 3v-2v=200 v=200mph

OpenStudy (anonymous):

Call wind speed w and speed of the plane in still air v. Head wind speed = v - w Tail wind speed = v + w Keep in mind that v > w. Otherwise, the plane will be flying "backward". Since the distance is the same in both trips, we can write (v - 40) x 3 = (v +40) x 2 3v - 120 = 2v + 80 v = 200 mph Note: The answer should be 200 mph, not 200 miles.